# The NHIM development scenario in an asymmetric binary system of dwarf galaxies

### v1

@@ 3.1 Development scenario of the most important periodic orbits
Now let us compare the actual Fig.2 with the Fig.11 in [21] for the corresponding symmetric case. Because of the symmetry in the old plot, we omitted the orbits coming out of $L_{3}$ and a large part of the orbits coming out of $P_{1}$. On the other hand, in the old plot, we included a few additional orbits which have turned out to be of little interest for the development scenario. In the new plot, these orbits are not included in order not to overload the plot. Therefore, to compare with the new asymmetric plot, we should imagine the old plot of the symmetric case augmented by the orbits coming out of $L_{3}$ and of $P_{1}$ and should delete the irrelevant orbits. Then we see a strong qualitative analogy between the two plots. There is only one essential difference which we point out now. In the new plot in Fig.2, we see two rather sharp corners in the red curves, one at $x\approx 0.3,E\approx-0.039$ and the other one at $x\approx 1.5,E\approx-0.035$. When the two masses become more similar and approach each other, then these two corners also approach each other in the middle and for equal masses, they touch and form a pitchfork bifurcation where lh1 absorbs a pair of other horizontal orbits coming from the potential minima. What we see in the asymmetric case in Fig.2 is a symmetry-broken pitchfork instead.
3.2 Development scenario of the internal dynamics of the NHIM1
(a) E=0.064E=-0.064
(b) E=0.046E=-0.046
(c) E=0.044E=-0.044
(d) E=0.042E=-0.042
(e) E=0.041E=-0.041
(f) E=0.04E=-0.04
Figure 3: Restricted Poincaré map of the NHIM1 for different energies
@@ 3.2 Development scenario of the internal dynamics of the NHIM1
We show the internal dynamics of NHIMs by the Poincaré map restricted to the NHIM surface with the intersection condition $z=0$ where the intersection orientation is irrelevant because of the symmetry under $z$ reflection. In the full 4-dimensional Poincaré map, the NHIM is a 2-dimensional curved surface embedded in the 4-dimensional domain of the map having coordinates $x,y,p_{x},p_{y}$. To display the restricted map graphically, we show the projection of the NHIM surface into some appropriate 2-dimensional coordinate surface. Here, the choice of the $y-p_{y}$ coordinate plane gave the best graphical results. Therefore, we will use this projection for all restricted Poincaré maps throughout the article.
The central and most important NHIM is the one over the Lagrange point $L_{1}$. Therefore, we start with the graphical presentation of the development scenario of this NHIM. It is created at the energy $E_{1}$. Very close to this energy, the internal dynamics of the NHIM is a two degrees of freedom (2-dof) oscillator, and accordingly, the Poincaré plot in Fig.3(a) for E=0.064E=-0.064 looks like that the one of a 2-dof oscillator. The central fixed point (a green dot in the figure) represents the vertical Lyapunov orbit lv1, and the boundary (shown as a solid red curve) is the horizontal Lyapunov orbit lh1, which lies completely in the horizontal plane. For such low energies, both of these Lyapunov orbits are tangentially elliptic. Now we increase the energy and show plots of the qualitatively changed NHIM1 appearance after each important bifurcation event, which is related to an important bifurcation event of the Lyapunov orbits, presented in the previous subsection.
@@
At an energy $E\approx-0.0638$, lh1 runs through its first pitchfork bifurcation, where it becomes tangentially hyperbolic and splits off a pair of tangentially elliptic tilted loop orbits. In the restricted map, these tilted loop orbits are represented by new fixed points which, at the bifurcation energy, come out of the boundary, i.e. out of lh1. In the figures, these tilted loop fixed points are represented by blue dots. The tilted loop orbits are surrounded by KAM islands, which grow larger with increasing energy. Fig.3(b) shows the situation for E=0.046E=-0.046. Here, the central island around lv1 is already rather small; the large majority of the area of the NHIM1 is occupied by the islands around the tilted loop orbits. In its first pitchfork bifurcation, lh1 has become tangentially unstable and has become part of a separatrix structure which divides the various islands. Strictly speaking, this separatrix structure is a fine chaos strip. However, it is so fine that in the figure it looks like lines. The tangential instability of lh1 is here still smaller than its normal instability. Accordingly, lh1 still belongs to the NHIM1.
@@
At an energy a little below E=0.045E=-0.045, lh1 runs through its second pitchfork bifurcation, where it splits off a new pair of tangentially hyperbolic tilted loop orbits, while lh1 itself returns to tangentially elliptic. The Fig.3(c) for the energy E=0.044E=-0.044 shows how these new tilted loop orbits run along the line $y=0$ closer to the central point for increasing energy. The 2 fixed points in the map, which correspond to these 2 new tilted loop orbits, are marked by magenta dots. The unstable manifolds of the new tangentially unstable tilted loop orbits create a fine chaos strip, which in the figure looks like a separatrix structure. This separatrix divides all the various disjoint regular islands. First, we have the very small remnant of the central island around lv1. Then we have the island pair around the tangentially elliptic tilted loop orbits. And last, we have a new regular region of KAM curves close to the boundary (which is still lh1), which has now become tangentially elliptic again.
@@
At the energy $E\approx-0.043$, the tangentially hyperbolic tilted loop orbits reach the central point and are absorbed by this central point in an inverse pitchfork bifurcation. Thereby, the central point turns tangentially hyperbolic and becomes the central point of the separatrix structure. The island around lv1 has been lost completely in this bifurcation. The Fig.3(d) shows the NHIM1 for E=0.042E=-0.042. We see how the separatrix structure has become a chaos strip of a size which is already clearly visible in the plot. In addition, we notice that the tilted loop islands start to disintegrate from the outside, i.e. from the neighbourhood of the chaos strip. Now we are in the energy region, which is most important for the whole development scenario. Therefore we show in Fig.3(e) the NHIM1 for the nearby energy E=0.041E=-0.041, even though no important bifurcation event happens during this further increase of the energy. Now the separatrix has already turned into a large-scale chaos region. It is important to emphasise that up to this energy, all elements in NHIM1 are still normally hyperbolic, i.e. in all points of the NHIM1, the normal instability is still larger than the tangential instability. This also holds for the large chaos region.
@@
At $E\approx-0.04035$, a very drastic event happens to the NHIM1. Here, the Lyapunov orbit lh1 loses its normal hyperbolicity. Accordingly, for higher values of the energy, lh1 no longer belongs to NHIM1, and it can no longer act as the boundary of NHIM1. At this energy, NHIM1 starts to decay. At first sight, the plot of NHIM1 for E=0.04E=-0.04 in Fig.3(f) does not look very different from the plot in Fig.3(e). Nevertheless, there are some important qualitative differences which we have to point out. As already mentioned, lh1 is no longer the boundary of NHIM1. Therefore, this orbit has been included in the plot as a broken line only. Now the large chaos strip no longer has any boundary curve on its outside, which prevents trajectories from diffusing chaotically to the outside in the long run and leaving the NHIM1 tangentially. Therefore, strictly speaking, the chaos strip is no longer invariant and accordingly no longer a part of the NHIM1. Trajectories only stay temporarily in the apparent chaos strip of the NHIM1. We can say that this apparent chaos strip only has a transient existence as an invariant surface. For more details on this type of transient behaviour on decaying NHIMs, see [23].
3.3 Development scenario of the NHIM3
@@
Next, let us have a look at NHIM3 over the Lagrange point $L_{3}$. This NHIM3 is created at the energy $E_{3}\approx-0.05921$. For energies a little larger than $E_{3}$, also this NHIM has an internal dynamics of a 2-dof oscillator, and accordingly, the reduced Poincaré map shows the corresponding pattern of concentric ellipses, where the central fixed point (again represented by a green dot in the figures) represents the vertical Lyapunov orbit lv3 and the boundary (again represented by a red solid curve in the figure) is the horizontal Lyapunov orbit lh3. It is not worth showing a plot for a very low energy.
(a) E=0.045E=-0.045
(b) E=0.043E=-0.043
(c) E=0.042E=-0.042
(d) E=0.041E=-0.041
(e) E=0.04E=-0.04
Figure 4: Restricted Poincaré map of the NHIM3 for different energies.
@@
At an energy $E\approx-0.055$, lh3 runs through a pitchfork bifurcation where it becomes tangentially hyperbolic and splits off a pair of tangentially elliptic tilted loop orbits, which in the reduced Poincaré map are the centres of islands and are again represented by blue dots in the figures. The Fig.4(a) shows the case of E=0.045E=-0.045 where the islands around the tilted loop orbits have already grown so large that they occupy approximately half of the area of NHIM3. Unlike lh1 and as we know already from subsection 3.1 and Fig.2, lh3 does not run through a further pitchfork bifurcation and remains tangentially hyperbolic as long as it belongs to NHIM3. Note that for the present energy, lh3 has a higher normal instability than its tangential instability; therefore, it is still normally hyperbolic and forms part of NHIM3. It is the outer part of a separatrix structure which it forms together with the separation line between the central island and the tilted loop islands. Strictly speaking, this separatrix structure is again a fine chaos strip; however, so fine that it appears as lines under the resolution of the graphics.
@@
We are already in the interesting energy region, therefore we increase $E$ in small steps and present in Fig.4(b) the reduced map of NHIM3 for E=0.043E=-0.043 even though the qualitative structure remains the same. We see that in comparison with Fig.4(a) many invariant lines from Fig.4(a) are broken into corresponding very fine secondary structures in Fig.4(b). This is a strong indication that we are already close to large-scale chaos and decay of the NHIM3.
@@
In Fig.4(c), we proceed to E=0.042E=-0.042 and see a very drastic change in comparison to Fig.4(b). As we already know from Fig.2, for energies a little smaller than -0.042, several important bifurcations of the important periodic orbits in NHIM3 occur. First, at $E\approx-0.0424$ lh3 loses its normal hyperbolicity and can no longer be a part of NHIM3. Here, the tangential instability has become larger than the normal instability. Second, at an energy extremely close to this event, at $E\approx-0.04239$, lh3 is destroyed in a saddle-centre bifurcation when it collides with another horizontal periodic orbit grown out of the potential minimum $P_{1}$. Accordingly, the outer boundary and the separatrix structure of the NHIM3 are lost. Third, at an energy $E\approx-0.04232$, the tilted loop orbits from NHIM3 collide with another pair of tilted loop orbits grown out of $P_{1}$ (see again Fig.2). Thereby, the tilted loop orbits together with their surrounding islands are lost from NHIM3. The only part of NHIM3 which persist this bifurcation process is the inner part of the central island around the vertical Lyapunov orbit lv3.
@@
In the region where the tilted loop islands have been destroyed, we find transient remnants of the NHIM3 in analogy to what we saw before in Fig.3(f) for NHIM1. In Fig.4(c), one transient trajectory in this decaying region is included by the orange crosses. We can imagine that in the decay of the 2-dimensional NHIM surface, lower-dimensional invariant subsets remain, and this collection of lower-dimensional invariant sets can hold general trajectories in their neighbourhood at least temporarily. Of course, these lower-dimensional unstable invariant sets also have stable and unstable manifolds of corresponding lower dimension, which can hold very particular incoming trajectories forever. But our numerical methods adapted to search for 2-dimensional NHIM surfaces and their 3-dimensional stable manifolds do not catch these lower-dimensional exceptional subsets. We only notice that general trajectories suffer some time delay in this region of the domain of the map.
@@
In Fig.4(d), we display the NHIM3 for E=0.041E=-0.041. With increasing energy, the size of the central island shrinks slowly. Here in the outer transient region, we encountered something instructive. We started a trajectory on the decaying NHIM surface at the initial coordinates $y=0$ and $p_{y}=-0.55$. This point happens to lie very close to the tangentially unstable central periodic point of a secondary 5:11 resonance. First, the iterates lie close to this periodic orbit before they deviate strongly for higher iterations and diffuse away. In total, we could follow this transient trajectory on the remnants of the NHIM3 for 125 iterations. This transient trajectory is again plotted by orange crosses. In Fig.4(e), we proceed to E=0.04E=-0.04. Compared to the previous part, there is little difference. Most interesting in this figure is the outer transient part. It is a single trajectory starting at $y=0$ and $p_{y}=-0.538$. The algorithm could find 81 iterations, the last one is the point near $y=4$, $p_{y}=-0.107$. We could not find a next iterate inside of the region belonging to the logical continuation of the NHIM3 to higher energies.
@@
For still higher energies, the central island of the NHIM3 shrinks slowly but remains. For $E\approx-0.028$ lv3 becomes tangentially unstable, but it remains normally hyperbolic even up to positive energies. The continuation of NHIM3 for such high energies is no longer relevant for the theme of the present article; therefore, we do not follow it any longer.
3.4 Development scenario of the NHIM2
Figure 5: Restricted Poincaré map of the NHIM2 at E=0.035E=-0.035.

### v2

@@ 3.1 Bifurcation scenario of the most important periodic orbits
To end this subsection, let us compare the actual Fig.2 with Fig.11 in [23] for the corresponding symmetric case. Because of the symmetry in the old plot, the orbits coming out of $L_{3}$ and a large part of the orbits coming out of $P_{1}$ are omitted. On the other hand, in the old plot are included a few additional orbits which have turned out to be of little interest for the bifurcation scenario. In the new plot, those irrelevant orbits are not included in order not to overload the plot. Therefore, to compare with the new asymmetric plot, we should imagine the old plot of the symmetric case augmented by the orbits coming out of $L_{3}$ and of $P_{1}$ and should delete the irrelevant orbits. Then we see a strong qualitative analogy between the two plots. There is only one essential difference which we point out now. In the new plot of Fig.2 we see two rather sharp corners in the red curves, one at $x\approx 0.3,E_{J}\approx-0.039$ and the other one at $x\approx 1.5,E_{J}\approx-0.035$. When the two masses become more similar and approach each other, then these two corners also approach each other in the middle and for equal masses, they touch and form a pitchfork bifurcation where $lh_{1}$ absorbs a pair of other horizontal orbits coming from the potential minima. What we see in the asymmetric case in Fig.2 is a symmetry-broken pitchfork instead.
3.2 Bifurcation scenario of the internal dynamics of the NHIM $\mathcal{M}^{1}_{E}$
(a) $E_{J}=-0.064$
(b) $E_{J}=-0.046$
(c) $E_{J}=-0.044$
(d) $E_{J}=-0.042$
(e) $E_{J}=-0.041$
(f) $E_{J}=-0.04$
Figure 3: Projected Poincaré map of the NHIM $\mathcal{M}^{1}_{E_{J}}$ and its remnant with transient trajectories for different values of the Jacobi constant $E_{J}$
@@ 3.2 Bifurcation scenario of the internal dynamics of the NHIM $\mathcal{M}^{1}_{E}$
We show the internal dynamics of NHIMs using the Poincaré map of the NHIM with the intersection condition $z=0$ where the intersection orientation is $p_{z}<0$. In the full 4-dimensional Poincaré map, the NHIM is a 2-dimensional curved surface embedded in the 4-dimensional domain of the map having coordinates $x,y,p_{x},p_{y}$. To display the Poincaré map graphically, we show the projection of the Poincaré map of the NHIM surface into some appropriate 2-dimensional surface. Here, the choice of the $yp_{y}$ canonical plane gave the best graphical results. Therefore, we will use this projection for all projected Poincaré maps throughout the article.
The central and most important NHIM is the one associated with the Lagrange point $L_{1}$, let us denote this NHIM as $\mathcal{M}^{1}_{E_{J}}$ for the value $E_{J}$ of the Jacobi constant. This NHIM is created at the Jacobi constant $E_{J1}$. Then, we start with the graphical presentation of the bifurcation scenario of this NHIM. Very close to this Jacobi constant $E_{J1}$, the internal dynamics of the NHIM $\mathcal{M}^{1}_{E_{J}}$ is a two degrees of freedom (2-dof) oscillator, and accordingly, the projected Poincaré map in Fig.3(a) for $E_{J}=-0.064$ looks like the one of a 2-dof oscillator. The red solid curve displays the horizontal Lyapunov orbit $lh_{1}$, which lies completely within the horizontal plane $z=0$. It marks the projection of the boundary of the NHIM $\mathcal{M}^{1}_{E_{J}}$ into the intersection plane $z=0$. When points on the projected Poincaré map of NHIM $\mathcal{M}^{1}_{E_{J}}$ approach this red boundary curve, then the value of $p_{z}$ of their corresponding trajectories in the moment of intersection approaches 0.
For low energies, both of the Lyapunov orbits are tangentially elliptic. Now we increase the Jacobi constant and show plots of the qualitatively changed NHIM $\mathcal{M}^{1}_{E_{J}}$ appearance after each important bifurcation event, which is related to an important bifurcation event of the Lyapunov orbits, presented in the previous subsection.
@@
At a Jacobi constant $E_{J}\approx-0.0638$, the periodic orbit $lh_{1}$ experiences its first pitchfork bifurcation, where it becomes tangentially hyperbolic and splits off a pair of tangentially elliptic tilted loop orbits. In the projected Poincaré map, these tilted loop orbits are represented by new fixed points which, at the bifurcation Jacobi constant, come out of the boundary, i.e. out of $lh_{1}$. In the figures, these tilted fixed points are represented by blue dots. The tilted loop orbits are surrounded by KAM tori, which grow larger with increasing Jacobi constant. Fig.3(b) shows the situation for $E_{J}=-0.046$. Here, the central KAM tori around $lv_{1}$ is already rather small; the large majority of the NHIM $\mathcal{M}^{1}_{E_{J}}$ is occupied by the KAM tori around the tilted loop orbits. In its first pitchfork bifurcation, $lh_{1}$ has become tangentially unstable and has become part of a separatrix structure which divides the various islands. Strictly speaking, this separatrix structure is a fine chaos strip. However, it is so fine that in the figure it looks like lines. The tangential instability of $lh_{1}$ is here still smaller than its normal instability. Accordingly, $lh_{1}$ still belongs to the NHIM $\mathcal{M}^{1}_{E_{J}}$.
@@
At a Jacobi constant a little below $E_{J}=-0.045$, the periodic orbit $lh_{1}$ experiences its second pitchfork bifurcation, where it splits off a new pair of tangentially hyperbolic tilted loop orbits, while $lh_{1}$ itself returns to tangentially elliptic. The Fig.3(c) for the Jacobi constant $E_{J}=-0.044$ shows how these new tilted loop orbits run along the line $y=0$ closer to the central point for increasing Jacobi constant. The 2 fixed points in the map, which correspond to these 2 new tilted loop orbits, are marked by magenta dots. The unstable manifolds of the new tangentially unstable tilted loop orbits create a fine chaos strip in the map, which, in the figure, looks like a separatrix structure. This separatrix divides all the various disjoint regular islands. First, we have the very small remnant of the central KAM tori around $lv_{1}$. Then we have the island pair around the tangentially elliptic tilted loop orbits. And last, In the projected Poincaré map, we have a new regular region of KAM curves close to the boundary (which is still $lh_{1}$), which has now become tangentially elliptic again.
@@
At the Jacobi constant $E_{J}\approx-0.043$, the projected Poincaré map of the tangentially hyperbolic tilted loop orbits reach the central point and are absorbed by this central point in an inverse pitchfork bifurcation. Thereby, the central point turns tangentially hyperbolic and becomes the central point of the separatrix structure. The island around $lv_{1}$ in the projected Poincaré map has been lost completely in this bifurcation. The Fig.3(d) shows the projected Poincaré map of NHIM $\mathcal{M}^{1}_{E_{J}}$ for $E_{J}=-0.042$. We see how the separatrix structure has become a chaos strip of a size which is already clearly visible in the plot. In addition, we notice that the tilted loop islands start to break from the outside, i.e. from the neighbourhood of the chaos strip. Now, the system is in the Jacobi constant region, which is most important for the whole bifurcation scenario. Therefore, we show in Fig.3(e) the projected Poincaré map of NHIM $\mathcal{M}^{1}_{E_{J}}$ for the nearby Jacobi constant $E_{J}=-0.041$, even though no important bifurcation event happens during this further increase of the Jacobi constant. Now the separatrix has already turned into a large-scale chaos region. It is important to emphasise that up to this Jacobi constant, the NHIM $\mathcal{M}^{1}_{E_{J}}$ has not yet started its breaking process, i.e. in all points of the NHIM $\mathcal{M}^{1}_{E_{J}}$, the normal instability is still larger than the tangential instability. This also holds for the large chaos region in the Poincaré map of NHIM $\mathcal{M}^{1}_{E_{J}}$.
@@
At $E_{J}\approx-0.04035$, a very drastic event happens to the NHIM $\mathcal{M}^{1}_{E_{J}}$. Here, the Lyapunov orbit $lh_{1}$ loses its normal hyperbolicity and the KAM closed invariant curves in its neighbourhood disappear. Accordingly, for higher values of the Jacobi constant, $lh_{1}$ no longer belongs to NHIM $\mathcal{M}^{1}_{E_{J}}$, and it can no longer act as the boundary of the projected Poincaré map of NHIM $\mathcal{M}^{1}_{E_{J}}$. At this Jacobi constant, NHIM $\mathcal{M}^{1}_{E_{J}}$ starts to break. At first sight, the plot of the projected Poincaré map of NHIM $\mathcal{M}^{1}_{E_{J}}$ for $E_{J}=-0.04$ in Fig.3(f) does not look very different from the plot in Fig.3(e). Nevertheless, there are some important qualitative differences which we have to point out. As already mentioned, $lh_{1}$ is no longer the boundary of the projected Poincaré map of the NHIM $\mathcal{M}^{1}_{E_{J}}$. Then, this orbit has been included in the plot as a broken line only. The large chaos strip that existed before became a region with transient chaos because there is no longer any KAM invariant torus which prevents trajectories from diffusing chaotically to the outside in the long run and leaving the NHIM $\mathcal{M}^{1}_{E_{J}}$ tangentially. Trajectories only stay temporarily in the transient chaos region around the NHIM $\mathcal{M}^{1}_{E_{J}}$. For more details on this type of transient behaviour on broken NHIMs, see [25].
3.3 Bifurcation scenario of the NHIM $\mathcal{M}^{3}_{E_{J}}$
@@
Next, let us have a look at NHIM $\mathcal{M}^{3}_{E_{J}}$ associated with the Lagrange point $L_{3}$. This NHIM $\mathcal{M}^{3}_{E_{J}}$ is created at the Jacobi constant $E_{J3}\approx-0.05921$. For energies a little larger than $E_{J3}$, also this NHIM has an internal dynamics of a 2-dof oscillator, and accordingly, the projected Poincaré map shows the corresponding pattern of concentric ellipses, where the central fixed point (again represented by a green dot in the figures) represents the vertical Lyapunov orbit $lv_{3}$ and the boundary (again represented by a red solid curve in the figure) is the horizontal Lyapunov orbit $lh_{2}$. It is not worth showing a plot for a very low Jacobi constant.
(a) $E_{J}=-0.045$
(b) $E_{J}=-0.043$
(c) $E_{J}=-0.042$
(d) $E_{J}=-0.041$
(e) $E_{J}=-0.04$
Figure 4: Projected Poincaré map of the NHIM $\mathcal{M}^{3}_{E_{J}}$ and its remnant with transient trajectories for different values of the Jacobi constant $E_{J}$
@@
At a Jacobi constant $E_{J}\approx-0.055$, $lh_{2}$ experiences a pitchfork bifurcation where it becomes tangentially hyperbolic and splits off a pair of tangentially elliptic tilted loop orbits, which in the projected Poincaré map are the centres of KAM islands and are again represented by blue dots in the figures. The Fig.4(a) shows the case of $E_{J}=-0.045$ where the islands around the tilted loop orbits have already grown so large that they occupy approximately half of the area of the projected Poincaré map of NHIM $\mathcal{M}^{3}_{E_{J}}$. Unlike $lh_{1}$ and as we know already from subsection 3.1 and Fig.2, $lh_{2}$ does not run through a further pitchfork bifurcation and remains tangentially hyperbolic as long as it belongs to NHIM $\mathcal{M}^{3}_{E_{J}}$. Note that for the present Jacobi constant, $lh_{2}$ has a higher normal instability than its tangential instability; therefore, it is still normally hyperbolic and forms part of NHIM $\mathcal{M}^{3}_{E_{J}}$. In the projected Poincaré map, it is the outer part of a separatrix structure which it forms together with the separation line between the central island and the tilted loop islands. Strictly speaking, this separatrix structure is again a fine chaos strip; however, so fine that it appears as lines under the resolution of the graphics.
@@
We are already in the interesting Jacobi constant region, therefore we increase $E_{J}$ in small steps and present in Fig.4(b) the projected map of NHIM $\mathcal{M}^{3}_{E_{J}}$ for $E_{J}=-0.043$ even though the qualitative structure remains the same. We see that in comparison with Fig.4(a) many invariant closed curves from Fig.4(a) are broken into corresponding very fine secondary structures in Fig.4(b). This is a strong indication that the system is already close to large-scale chaos and break of the NHIM $\mathcal{M}^{3}_{E_{J}}$.
@@
In Fig.4(c), we proceed to $E_{J}=-0.042$ and see a very drastic change in comparison to Fig.4(b). As we already know from Fig.2, for a Jacobi constant a little smaller than $E_{J}=-0.042$, several important bifurcations of the important periodic orbits in NHIM $\mathcal{M}^{3}_{E_{J}}$ occur. First, at $E_{J}\approx-0.0424$ $lh_{2}$ loses its normal hyperbolicity and can no longer be a part of NHIM $\mathcal{M}^{3}_{E_{J}}$. Here, the tangential instability has become larger than the normal instability. Second, at a Jacobi constant extremely close to this event, at $E_{J}\approx-0.04239$, $lh_{2}$ is destroyed in a saddle-centre bifurcation when it collides with another horizontal periodic orbit grown out of the potential minimum $P_{1}$. Accordingly, the outer boundary and the separatrix structure of the NHIM $\mathcal{M}^{3}_{E_{J}}$ are lost. Third, at a Jacobi constant $E_{J}\approx-0.04232$, the tilted loop orbits from NHIM $\mathcal{M}^{3}_{E_{J}}$ collide with another pair of tilted loop orbits grown out of $P_{1}$ (see Fig.2). Thereby, the tilted loop orbits together with their surrounding KAM tori are lost from NHIM $\mathcal{M}^{3}_{E_{J}}$. The only part of NHIM $\mathcal{M}^{3}_{E_{J}}$ that persists this bifurcation process is the inner part of the central KAM tori around the vertical Lyapunov orbit $lv_{3}$.
@@
In the transient region where the tilted loop orbits have been destroyed, we find transient remnants of NHIM $\mathcal{M}^{3}_{E_{J}}$ in analogy to what we saw before in Fig.3(f) for NHIM $\mathcal{M}^{1}_{E_{J}}$. In Fig.4(c), the projection of the Poincaré map of one transient trajectory in the region is included by the orange crosses. We can imagine that in the break of the 2-dimensional projected Poincaré map of the NHIM, lower-dimensional invariant subsets remain, and this collection of lower-dimensional invariant sets can hold general trajectories in their neighbourhood at least temporarily. Of course, these lower-dimensional unstable invariant sets also have stable and unstable manifolds of corresponding lower dimension, which can hold very particular incoming trajectories forever. But our numerical methods adapted to search for a 2-dimensional projected Poincaré map of the NHIM and their 3-dimensional stable manifolds do not catch these lower-dimensional exceptional subsets. We only notice that general trajectories experience some time delay in this region of the domain of the map.
@@
In Fig.4(d), we display the projection of Poincaré map of NHIM $\mathcal{M}^{3}_{E_{J}}$ for $E_{J}=-0.041$. With increasing Jacobi constant, the size of the central island shrinks slowly. Here in the outer transient region, we encountered something instructive. We started a trajectory in the transient region close to the projected Poincaré map of NHIM $\mathcal{M}^{3}_{E_{J}}$ at the initial coordinates $y=0$ and $p_{y}=-0.55$. This point happens to lie very close to the tangentially unstable central periodic point of a secondary 5:11 resonance. First, the iterates lie close to this periodic orbit before they deviate strongly for higher iterations and diffuse away. In total, we could follow this transient trajectory on the remnants of the NHIM $\mathcal{M}^{3}_{E_{J}}$ for 125 iterations of the Poincaré map. The map of this transient trajectory is again plotted by orange crosses. In Fig.4(e), we proceed to $E_{J}=-0.04$. Compared to the previous part, there is little difference. Most interesting in this figure is the outer transient part. It is a single trajectory starting at $y=0$ and $p_{y}=-0.538$. The algorithm could find 81 iterations of the Poincaré map, the last one is the point near $y=4$, $p_{y}=-0.107$. We could not find a next iterate inside the region surrounding the NHIM $\mathcal{M}^{3}_{E_{J}}$ to higher energies.
@@
For still higher energies, the central island of the NHIM $\mathcal{M}^{3}_{E_{J}}$ shrinks slowly but remains. For $E_{J}\approx-0.028$, $lv_{3}$ becomes tangentially unstable, but it remains normally hyperbolic even up to positive energies. The continuation of NHIM $\mathcal{M}^{3}_{E_{J}}$ for such high energies is no longer relevant for the theme of the present article; therefore, we do not follow it any longer.
3.4 Bifurcation scenario of the NHIM $\mathcal{M}^{2}_{E_{J}}$
Figure 5: Projected Poincaré map of the NHIM $\mathcal{M}^{2}_{E_{J}}$ for Jacobi constant $E_{J}=-0.035$.